21. y + 3y + 3y + y = 0 22. y – 6y + 12y – 8y = 0 23. y(a) + y + y =0 24. y(4) – 2y +y=0 In Problems 1-14 find the geCan you please show number #25, #27 (Please make work readable)
Answer #1

We will find the characteristic equation and its roots. Then we find the general solution using these roots. The detailed sol

25 The given differential equation is - 16.04 + 24 d y 7 + 99 = 0 The characteristic equation is 16m² + 24 m² +9=0 (4 m+ 3) == (x+1)} (x-1)} (x+5) = 0 x=-5 ou X = -1 (twice) or x=1 (twice) : The solution is y(x) = Geox + (6 + GXDē + (cy + GgXJL* whe